jlaustsen:

Find the exponential function that satisfies the given conditions: Initial value = 34, increasing at a rate of 7% per year

Laylalyssa:

Welcome to QC! ok so, f(x)=33*(1.07)^x f(x) will be ur exponential growth function showing growth after x years.u have to find the exponential function that works with the given conditions: Initial value = 33, increasing at a rate of 7% per year. bc an exponential growth function is in form: y=a*(1+r),where a= initial value of function and r = growth rate in decimal form. Given: a=33 r=7%. sooo convert ur given rate in decimal form. 7% = 7/100 =0.07- this is basic math(shouldn't even be considered math) then u have to substitute our given values in exponential function form: f(x)=33*(1+0.07)^x f(x)=33*(1.07)^x and nowwwww, the exponential function that satisfies our given conditions will be . f(x)=33*(1.07)*^x